# BEAST: Efficient Tokenization of B-Splines Encoded Action Sequences for Imitation Learning

### v2

@@ Abstract
We present the B-spline Encoded Action Sequence Tokenizer (BEAST), a novel action tokenizer that encodes action sequences into compact discrete or continuous tokens using B-spline. In contrast to existing action tokenizers based on vector quantization or byte pair encoding, BEAST requires no separate tokenizer training and consistently produces tokens of uniform length, enabling fast action sequence generation via parallel decoding. Leveraging our B-spline formulation, BEAST inherently ensures generating smooth trajectories without discontinuities between adjacent segments. We extensively evaluate BEAST by integrating it with three distinct model architectures: a Variational Autoencoder (VAE) with continuous tokens, a decoder-only Transformer with discrete tokens, and Florence-2, a Vision-Language Model with an encoder-decoder architecture, demonstrating BEAST’s compatibility and scalability with large pretrained models. We evaluate BEAST across three established benchmarks consisting of 166 simulated tasks and on three distinct robot settings with a total of 8 real-world tasks. Experimental results demonstrate that BEAST (i) significantly reduces both training and inference computational costs, and (ii) consistently generates smooth, high-frequency control signals suitable for continuous control tasks while (iii) reliably achieves competitive task success rates compared to state-of-the-art methods. Videos and code are available at the project page.
@@ 1 Introduction
Imitation learning has emerged as a powerful paradigm for training robots to perform complex tasks by learning from human demonstrations [1]. Early works [2, 3] in this field primarily focused on predicting single-step actions based on the current observation. However, recent research [4] highlights the importance of learning action sequences to capture the temporal coherence inherent in human demonstrations. Moreover, by modeling action sequences, we can reduce compounding errors [5] and create task demonstrations that more closely align with human methods [6]. Given the success of autoregressive next-token prediction models in natural language processing and other domains [7, 8, 9], it is compelling to explore similar techniques for modeling action sequences, leveraging their ability to predict and generate coherent sequences effectively.
@@
To address these challenges, we propose the B-spline Encoded Action Sequence Tokenizer (BEAST), a novel tokenizer that represents continuous action sequences using B-splines [19]. BEAST offers versatility, allowing for effective integration with both discrete and continuous tokens. Different from tokenizers based on the vector quantization [13, 14, 15], it does not require additional tokenizer training. BEAST compresses action trajectories into fixed-length token sequences enabling efficient parallel decoding for faster token generation, requiring 48×4-8$\times$4 - 8 × fewer tokens than binning-based tokenization. By using B-spline encoded control points as discrete tokens, BEAST ensures the generation of smooth action chunks, as well as the continuous connection between consecutive chunks.
@@ 2 Related Work
Action Representations for Imitation Learning. Prior work has explored various action representations for policy learning. The most common approach is to directly predict low-level actions, such as joint positions or end-effector displacements, using a supervised learning objective [3, 2, 20]. While simple, these approaches cannot tackle the multimodality present in human behavior. To address these limitations, ACT [4] introduces an Action Chunking Transformer trained as a conditional Variational Autoencoder (CVAE), which models multimodal behavior via a learned latent space. Instead of predicting single actions, ACT generates entire action chunks in a single inference step. These chunks are short sequences of actions, which reduces covariate shift and improves performance. Another line of work focuses on generating action sequences with diffusion models. Diffusion Policies model complex, multimodal behaviors by iteratively denoising from Gaussian noise to generate action sequences [5, 21, 22, 23]. While effective, these methods require multiple denoising steps per sequence, making inference comparatively expensive. In contrast, BEAST compresses full action sequences into compact control-point representations using B-spline approximation. This significantly reduces the number of predictions needed to model temporally extended behaviors. As a result, it enables efficient action chunking with smooth transitions, combining the representational benefits of ACT and diffusion policies with the speed and simplicity of tokenized inference.
@@ 3 Preliminaries
[…]
where ${\bm{c}}$ are $N$ control points and $u$ is a continuous parameter, often interpreted as normalized time. The basis functions $\bm{\Phi^{P}}(u)=[\Phi^{P}_{0}(u),..,\Phi^{P}_{N-1}(u)]$ are $N$ polynomial basis functions of $P$-th degree. These basis functions are defined over $M$ intervals determined by $M+1$ knots in a vector $[k_{0},...,k_{M}]$, and it satisfies $M=N+P$ [30]. Typically, the knot vector is normalized such that $k_{0}=0$ and $k_{M}=1$. The basis functions $\Phi^{P}_{n}(u)$ are recursively computed using the Cox–de Boor recursion [31]. We denote all recursive degrees [^1] The B-Spline degree $P$ differs from the recursive degree $q$. Trajectories are represented by basis functions of degree $P$, while lower recursive degree $q$ serve as intermediate representations in the recursive process. as $q=0:P$. For $q=0$, the basis functions are defined as piecewise constant and recursively using the $(q-1)$-th degree basis for $q>0$ with
[…]
@@ 4.1 Action Sequence Tokenization with B-Spline Tokenizer
Following prior works in action tokenization [10, 16], we first normalize the input actions such that the 1st and 99th quantile value of each action dimension in the dataset maps to the range of $[-1,1]$. Using quantiles makes the normalization robust against outlier data points.
@@
${\bm{c}}=\arg\min_{{\bm{c}}}||y_{1:T}-a_{1:T}||^{2}_{2}=\arg\min_{\bm{c}}||\bm{\Phi}^{P}(u){\bm{c}}-a_{1:T}||^{2}_{2},$ (4)
where $\bm{\Phi}^{P}(u)=[\Phi^{P}_{1}(u),\Phi^{P}_{2}(u),...,\Phi^{P}_{N}(u)]^{\top}$ represents precomputed B-spline basis functions defined over interval $u\in[0,1]$. Ridge regression estimates the control points in closed form, ${\bm{c}}=[c_{0},c_{1},...,c_{N-1}]=(\bm{\Phi}^{\top}\bm{\Phi}+\lambda\bm{I})^{-1}\bm{\Phi}^{\top}a_{1:T}$, with $\lambda$ acting as a regularization parameter. This efficient computation typically introduces only a small overhead, typically $3$ to $5$ milliseconds per batch. For a high-dimensional action sequence, i.e. $D\!>\!1$, each DoF is encoded independently into ${\bm{c}}_{d}$, resulting in a matrix ${\bm{C}}$ of shape $D\!\times\!N$, that stacks each DoF’s control points, ${\bm{C}}\!=\![{\bm{c}}_{1},{\bm{c}}_{2},...,{\bm{c}}_{D}]^{\!\top}.\,\,$
@@ 4.2 Enforcing Smooth Transition with Clamped B-Spline
In contrast, BEAST employs clamped B-Spline to ensure smooth transitions between consecutive action chunks. As introduced in Section 3, clamped B-Spline is a specialized variant of B-Spline that guarantees to start from the first control point and end at the last control point, which is utilized to generate seamlessly connected action sequences, as illustrated in Figure 1. To ensure smooth transitions, we directly set the first control point ${\bm{c}}_{0}$ to the last action of the previous sequence. We then compute the residual trajectory $\hat{{\bm{a}}}$ by subtracting the contribution of the first basis function: $\hat{{\bm{a}}}={\bm{a}}-c_{0}\Phi_{0}^{P}$. The remaining control points $\hat{{\bm{c}}}=[c_{1},c_{2},...c_{N-1}]$ are determined by solving the linear regression problem similar to equation 4: $\arg\min_{\bm{c}}||\hat{\Phi}^{P}(u)\hat{{\bm{c}}}-\hat{{\bm{a}}}||_{2}$. Through this approach, BEAST consistently generates action sequences with mathematically guaranteed smooth transitions between chunks. This will be further discussed in our toy task experiment in Section 5.1.
@@ 4.3 Combining BEAST tokens with different architectures
Continuous Tokens. We also explore the performance of combining BEAST with ACT [4]. ACT uses a conditional VAE (CVAE) with a Transformer Encoder-Decoder to predict a sequence of actions. We predict $N$ BEAST continuous tokens, where each token has the dimension of $D$, instead of action sequences, this design choice keeps the temporal order inherent in the trajectory segments. Using the BEAST tokenization, we reduce the length of predicted token sequence by $6.67$ times (from 100 to 15) without sacrificing the task performance. In addition, our method enables smooth trajectories without requiring temporal aggregation.
@@ 5 Experiments
We conducted extensive evaluations in both simulated and real-world settings, targeting answering five key research questions (RQs): 1) What advantages does BEAST offer over commonly used binning-based tokenizers? 2) How does BEAST contribute to the performance on imitation learning benchmarks? 3) How does BEAST affect the training and inference efficiency? 4) Does BEAST generalize to real-world scenarios? 5) How do the design choices affect the performance of BEAST? BEAST is integrated into two different architectures: We combine BEAST and Florence-2 [32] and term this VLA variant as BEAST-F. In contrast to many baselines, we test BEAST-F without second-stage pretraining on large-scale robot datasets. We additionally apply BEAST to a small decoder-only transformer model (BEAST-D) and on top of a vanilla ACT [4] (BEAST-ACT).
@@ 5.3 Advantages in Training and Inference Speed
Next, we verify the inference and training efficiency of BEAST to answer RQ3. Specifically, we consider the VLA variant BEAST-F and compare it against several recent VLAs [10, 43, 16], as well as a standard CNN-based Diffusion Policy[5]. We measure the inference efficiency on an RTX 4090 GPU. As shown in Table  3, BEAST-F demonstrates clear computational advantages. It achieves a throughput of 617.3 Hz (e.g., generates approximately 617 actions per second), which is 2.14×2.14$\times$2.14 × faster than $\pi_{0}$, 4.72×4.72$\times$4.72 × faster than Diffusion Policy, and 101.4×101.4$\times$101.4 × faster than OpenVLA. In addition, BEAST-F achieves the lowest latency at just $19$ milliseconds, where latency refers to the time taken to generate one action chunk. These gains are due to the parallel decoding, which enables generating the action sequence in a single forward pass.
[…]
@@ 5.4 Real-World Evaluation with 3 Different Robot Setups
To answer RQ4, we assess the effectiveness of BEAST across diverse real-world scenarios with varying data collection frequencies. We evaluate BEAST on 8 challenging manipulation tasks across 3 different experimental setups: 1) Franka Challenge: Four tabletop manipulation tasks (Towel Fold, Sweep, Mixer, Pour) using a joint position-controlled Franka robot with data collected at 20Hz, 2) Real Kitchen: Three manipulation tasks on a toy kitchen setup (Move Banana, Open Oven, Move Pot) with data collected at 35Hz, 3) Bi-manual ALOHA: A cube transfer task using a bi-manual ALOHA robot with data recorded at 60Hz. For each task in the Franka Challenge and Franka Kitchen setups, we conduct 10 evaluation runs per method, while for the ALOHA cube transfer task, we performed 30 runs. The average success rate for each task is reported in Figure 8. For tasks comprising multiple stages, we track intermediate milestones to better evaluate the completion of each sub-task. Appendix D provides a detailed description of all setups and tasks. We compare BEAST against $\pi_{0}$[43], $\pi_{0}$-FAST[16], and ACT [4]. We finetune $\pi_{0}$ and $\pi_{0}$-FAST from the official pretrained checkpoints for an additional 60k and 40k steps, respectively. For each method, we train one multitask model for all four tasks, Real Franka tasks, and another for the Real Kitchen tasks. The results demonstrate that BEAST-F achieves $52.86\%$ success rate and BEAST-D achieves $76.57\%$. In contrast $\pi_{0}$ achieves $53.43\%$ and FAST only $28.5\%$. Interestingly, the smaller model (BEAST-D) outperforms all the VLAs, including the Florence variant with BEAST. We attribute this effect to the relatively small real-world dataset of only 50 demonstrations for each task. For the Aloha Cube Transfer task, we compare BEAST-ACT against the base ACT that directly predicts action sequences in the joint space. BEAST-ACT achieves $70\%$ success, which is $21\%$ higher than the base ACT.
@@
Figure 8: Experimental Results on Real-World Robot Tasks. This figure shows the average task success rate across eight real-world tasks. Each task and method was evaluated over 10 runs (30 runs for Cube Transfer). Success rates are measured at the sub-task level. Detailed descriptions of all sub-tasks are provided in Appendix D. BEAST variants achieve strong performance in real world.
@@ 5.5 Ablation Studies
BEAST-F [10] 4.43
BEAST-F [5] 3.88
BEAST-F [15] 4.14
BEAST-F [20] 2.71
@@
Table 4: Average Sequence Lengths for BEAST-F Ablations on CALVIN ABC.
@@
Choice of Number of Basis Functions. Next, we evaluate how the number of basis functions affects the policy performance. We evaluate using $N=[5,10,15,20]$ basis functions to model action chunks of $20$ steps, denoted as BEAST-F [N] in Table 4. Fewer basis functions lead to fewer tokens for prediction, but it also reduces the expressiveness of the B-Spline representation. On the contrary, more basis functions increase representational power but reduce compression, which can also negatively influence the performance.
Scaling with Model Size. Finally, we assess the impact of model size on task performance. We compare BEAST-F, which uses Florence-2-large ($0.77$B parameters), with BEAST-SF, a smaller variant based on Florence-2-base ($0.23$B parameters). The larger model achieves an $11.3\%$ improvement in average sequence length, demonstrating that BEAST benefits from increased model capacity. This result highlights its potential as a scalable building block for larger VLAs.

### v3

@@ Abstract
We present the B-spline Encoded Action Sequence Tokenizer (BEAST), a novel action tokenizer that encodes action sequences into compact discrete or continuous tokens using B-spline. In contrast to existing action tokenizers based on vector quantization or byte pair encoding, BEAST requires no separate tokenizer training and consistently produces tokens of uniform length, enabling fast action sequence generation via parallel decoding. Leveraging our B-spline formulation, BEAST inherently ensures generating smooth trajectories without discontinuities between adjacent segments. We extensively evaluate BEAST by integrating it with three distinct model architectures: a Variational Autoencoder (VAE) with continuous tokens, a decoder-only Transformer with discrete tokens, and Florence-2, a Vision-Language Model with an encoder-decoder architecture, demonstrating BEAST’s compatibility and scalability with large pretrained models. We evaluate BEAST across three established benchmarks consisting of 166 simulated tasks and on three distinct robot settings with a total of 8 real-world tasks. Experimental results demonstrate that BEAST (i) significantly reduces both training and inference computational costs, and (ii) consistently generates smooth, high-frequency control signals suitable for continuous control tasks while (iii) reliably achieves competitive task success rates compared to state-of-the-art methods. Videos and code are available at https://intuitive-robots.github.io/beast_website/.
@@ 1 Introduction
Imitation learning has emerged as a powerful paradigm for training robots to perform complex tasks by learning from human demonstrations [1, 2, 3]. Early works [4, 5] in this field primarily focused on predicting single-step actions based on the current observation. However, recent research [6] highlights the importance of learning action sequences to capture the temporal coherence inherent in human demonstrations. Moreover, by modeling action sequences, we can reduce compounding errors [7] and create task demonstrations that more closely align with human methods [8]. Given the success of autoregressive next-token prediction models in natural language processing and other domains [9, 10, 11], it is compelling to explore similar techniques for modeling action sequences, leveraging their ability to predict and generate coherent sequences effectively.
@@
To address these challenges, we propose the B-spline Encoded Action Sequence Tokenizer (BEAST), a novel tokenizer that represents continuous action sequences using B-splines [21]. BEAST offers versatility, allowing for effective integration with both discrete and continuous tokens. Different from tokenizers based on the vector quantization [15, 16, 17], it does not require additional tokenizer training. BEAST compresses action trajectories into fixed-length token sequences enabling efficient parallel decoding for faster token generation, requiring $4-8\times$ fewer tokens than binning-based tokenization. By using B-spline encoded control points as discrete tokens, BEAST ensures the generation of smooth action chunks, as well as the continuous connection between consecutive chunks.
@@ 2 Related Work
Prior work has explored various action representations for policy learning. The most common approach is to directly predict low-level actions, such as joint positions or end-effector displacements, using a supervised learning objective [5, 4, 22]. While simple, these approaches cannot tackle the multimodality present in human behavior [23].
To address these limitations, ACT [6] introduces an Action Chunking Transformer trained as a conditional Variational Autoencoder (CVAE), which models multimodal behavior via a learned latent space. Instead of predicting single actions, ACT generates entire action chunks in a single inference step. These chunks are short sequences of actions, which reduces covariate shift and improves performance. Another line of work focuses on generating action sequences with diffusion models. Diffusion Policies model complex, multimodal behaviors by iteratively denoising from Gaussian noise to generate action sequences [7, 24, 25, 26, 27, 28]. While effective, these methods require multiple denoising steps per sequence, making inference comparatively expensive. In contrast, BEAST compresses full action sequences into compact control-point representations using B-spline approximation. This significantly reduces the number of predictions needed to model temporally extended behaviors. As a result, it enables efficient action chunking with smooth transitions, combining the representational benefits of ACT and diffusion policies with the speed and simplicity of tokenized inference.
A prominent research direction in action sequence representation is Movement Primitives (MPs) [29, 30, 31, 32]. Dynamic Movement Primitives (DMPs) [29, 31] model trajectories using a second-order dynamical system with a goal attractor, ensuring smooth transitions in position and velocity. However, DMPs are limited in representing higher-order continuity. To address these limitations, recent works [32, 33] have explored using B-splines as an alternative representation for MPs. Despite their effectiveness, existing works have primarily applied MPs within reinforcement learning (RL) frameworks [34, 35, 36, 37, 38, 33] and only utilize them as continuous action representations.
@@ 3 Preliminaries
$\textbf{1-DoF B-Spline:}\quad\quad y(u)=\sum_{n=0}^{N-1}\Phi_{n}^{P}(u)~c_{n}=\bm{\Phi}^{P}(u)~{\bm{c}},\quad 0\leq P<N,\quad u\in[k_{0},k_{M}],$ (1)
where $\bm{c}$} are $N$ control points and $u$ is a continuous parameter, often interpreted as normalized time. The basis functions $\bm{\Phi^{P}}(u)=[\Phi^{P}_{0}(u),..,\Phi^{P}_{N-1}(u)]$ are $N$ polynomial basis functions of $P$-th degree. These basis functions are defined over $M$ intervals determined by $M+1$ knots in a vector $k_{0},...,k_{M}]$, and it satisfies $M=N+P$ [45]. Typically, the knot vector is normalized such that $k_{0}=0$ and $k_{M}=1$. The basis functions $\Phi^{P}_{n}(u)$ are recursively computed using the Cox–de Boor recursion [46]. We denote all recursive degrees [^1] The B-Spline degree $P$ differs from the recursive degree $q$. Trajectories are represented by basis functions of degree $P$, while lower recursive degree $q$ serve as intermediate representations in the recursive process. as $q=0:P$. For $q=0$, the basis functions are defined as piecewise constant and recursively using the (q1)(q-1)-th degree basis for $q>0$ with
$\textbf{piecewise constant:}\qquad\Phi_{n}^{0}(u)$ $=\begin{cases}1&\text{if }k_{n}\leq u<k_{n+1},\\ 0&\text{otherwise}.\end{cases}\textnormal{and}$ (2)
$\textbf{recursive:}\qquad\Phi_{n}^{q}(u)$ $=k_{n}^{q-1}\Phi_{n}^{q-1}(u)+(1-k_{n+1}^{q-1})\Phi_{n+1}^{q-1}(u),$ (3)
@@ 4.1 Action Sequence Tokenization with B-Spline Tokenizer
Following prior works in action tokenization [12, 18], we first normalize the input actions such that the 1st and 99th quantile value of each action dimension in the dataset maps to the range of [1,1][-1,1]. Using quantiles makes the normalization robust against outlier data points.
@@
[…]
where $\bm{\Phi}^{P}(u)=[\Phi^{P}_{1}(u),\Phi^{P}_{2}(u),...,\Phi^{P}_{N}(u)]^{\top}$ represents precomputed B-spline basis functions defined over interval $u\in[0,1]$. Ridge regression estimates the control points in closed form, ${\bm{c}}=[c_{0},c_{1},...,c_{N-1}]=(\bm{\Phi}^{\top}\bm{\Phi}+\lambda\bm{I})^{-1}\bm{\Phi}^{\top}a_{1:T}$, with $\lambda$ acting as a regularization parameter. This efficient computation typically introduces only a small overhead, typically $3$ to $5$ milliseconds per batch. For a high-dimensional action sequence, i.e. D>1D\!>\!1, each DoF is encoded independently into ${\bm{c}}_{d}$, resulting in a matrix $\bm{C}$} of shape $D\!\times\!N$, that stacks each DoF’s control points, ${\bm{C}}\!=\![{\bm{c}}_{1},{\bm{c}}_{2},...,{\bm{c}}_{D}]^{\!\top}.\,\,$
@@ 4.2 Enforcing Smooth Transition with Clamped B-Spline
[…]
@@ 4.3 Combining BEAST tokens with different architectures
Continuous Tokens. We also explore the performance of combining BEAST with ACT [6]. ACT uses a conditional VAE (CVAE) with a Transformer Encoder-Decoder to predict a sequence of actions. We predict $N$ BEAST continuous tokens (normalized B-Spline control points without quantization), where each token has the dimension of $D$, instead of action sequences, this design choice keeps the temporal order inherent in the trajectory segments. Using BEAST, we reduce the length of predicted token sequence by $6.67$ times (from 100 to 15) without sacrificing the task performance. In addition, our method enables smooth trajectories without requiring temporal aggregation.
@@ 5 Experiments
We conducted extensive evaluations in both simulated and real-world settings, targeting answering five key research questions (RQs): 1) What advantages does BEAST offer over commonly used binning-based tokenizers? 2) How does BEAST contribute to the performance on imitation learning benchmarks? 3) How does BEAST affect the training and inference efficiency? 4) Does BEAST generalize to real-world scenarios? 5) How do the design choices affect the performance of BEAST? BEAST is integrated into three different architectures: First, we combine BEAST and Florence-2 [47] and term this VLA variant as BEAST-F. Second, we integrate BEAST into a small decoder-only transformer trained from scratch, referred to as BEAST-D. Finally, we employ continuous BEAST tokens on top of a vanilla ACT[6], resulting in (BEAST-ACT). A detailed description of each architecture is provided in Appendix B. In contrast to many baselines, we test BEAST-F without second-stage pretraining on large-scale robot datasets.
@@ 5.3 Advantages in Training and Inference Speed
Next, we verify the inference and training efficiency of BEAST to answer RQ3. Specifically, we consider the VLA variant BEAST-F and compare it against several recent VLAs [12, 58, 18], as well as a standard CNN-based Diffusion Policy[7]. We measure the inference efficiency on an RTX 4090 GPU. As shown in Table  3, BEAST-F demonstrates clear computational advantages. It achieves a throughput of 617.3 Hz (e.g., generates approximately 617 actions per second), which is $2.14\times$ faster than $\pi_{0}$, $4.72\times$ faster than Diffusion Policy, and $101.4\times$ faster than OpenVLA. In addition, BEAST-F achieves the lowest latency at just $19$ milliseconds, where latency refers to the time taken to generate one action chunk. These gains are due to the parallel decoding, which enables generating the action sequence in a single forward pass.
10k20k30k00.20.20.40.40.60.60.80.811Training StepsSuccess RateBEAST-$\pi_{0}F$AST
@@
5.4 Tokenizer Comparison under a Unified Backbone
LIBERO-LONGCALVIN ABC-D0505010010030302828848446468484717186868888Success Rate (%)Binning-ARFAST-ARBEAST-ARBEAST-PD
Figure 8: Tokenizer Comparison.
[…]
@@ 5.5 Real-World Evaluation with 3 Different Robot Setups
To answer RQ4, we assess the effectiveness of BEAST across diverse real-world scenarios with varying data collection frequencies. We evaluate BEAST on 8 challenging manipulation tasks across 3 different experimental setups: 1) Franka Challenge: Four tabletop manipulation tasks (Towel Fold, Sweep, Mixer, Pour) using a joint position-controlled Franka robot with data collected at 20Hz, 2) Real Kitchen: Three manipulation tasks on a toy kitchen setup (Move Banana, Open Oven, Move Pot) with data collected at 35Hz, 3) Bi-manual ALOHA: A cube transfer task using a bi-manual ALOHA robot with data recorded at 60Hz. For each task in the Franka Challenge and Franka Kitchen setups, we conduct 10 evaluation runs per method, while for the ALOHA cube transfer task, we performed 30 runs. The average success rate for each task is reported in Figure 9. For tasks comprising multiple stages, we track intermediate milestones to better evaluate the completion of each sub-task. Appendix F provides a detailed description of all setups and tasks. We compare BEAST against $\pi_{0}$[58], $\pi_{0}$-FAST[18], and ACT [6]. We finetune $\pi_{0}$ and $\pi_{0}$-FAST from the official pretrained checkpoints for an additional 60k and 40k steps, respectively. For each method, we train one multitask model for all four tasks, Real Franka tasks, and another for the Real Kitchen tasks. The results demonstrate that BEAST-F achieves $52.86\%$ success rate and BEAST-D achieves $76.57\%$. In contrast $\pi_{0}$ achieves $53.43\%$ and FAST only $28.5\%$. Interestingly, the smaller model (BEAST-D) outperforms all the VLAs, including the Florence variant with BEAST. We attribute this effect to the relatively small real-world dataset of only 50 demonstrations for each task. For the Aloha Cube Transfer task, we compare BEAST-ACT against the base ACT that directly predicts action sequences in the joint space. BEAST-ACT achieves $70\%$ success, which is $21\%$ higher than the base ACT.
@@
Figure 9: Experimental Results on Real-World Robot Tasks. This figure shows the average task success rate across eight real-world tasks. Each task and method was evaluated over 10 runs (30 runs for Cube Transfer). Success rates are measured at the sub-task level. Detailed descriptions of all sub-tasks are provided in Appendix F. BEAST variants achieve strong performance in real world.
@@ 5.6 Ablation Studies
BEAST-F (10) 4.43
BEAST-F (5) 3.88
BEAST-F (15) 4.20
BEAST-F (20) 4.32
BEAST-F (25) 4.23
@@
Table 4: Average Sequence Lengths for BEAST-F Ablation.
@@
Choice of Number of Basis Functions. Next, we evaluate how the number of basis functions affects the policy performance. We evaluate using $N=[5,10,15,20,25]$ basis functions to model action chunks of $20$ steps, denoted as BEAST-F [N] in Table 4. Fewer basis functions lead to fewer tokens for prediction, but it also reduces the expressiveness of the B-Spline representation. On the contrary, more basis functions increase representational power but reduce compression, which can also negatively influence the performance.
Scaling with Model Size. Finally, we assess the impact of model size on task performance. We compare BEAST-F, which uses Florence-2-large (0.770.77B parameters), with BEAST-SF, a smaller variant based on Florence-2-base (0.230.23B parameters). The larger model achieves an $11.3\%$ improvement in average sequence length, demonstrating that BEAST benefits from increased model capacity. This result highlights its potential as a scalable building block for larger VLAs.